OpenStudy (anonymous):

8x^3-27 how do you do it

OpenStudy (anonymous):

(2 x-3) (4 x^2+6 x+9)

OpenStudy (anonymous):

\[(2x-3)^{3}\]

OpenStudy (anonymous):

This is a standard formula, difference of cubes a^3-b^3 = (a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

If what you're being asked to do is factor, then that's the answer. If you're being asked to solve it I can give you another method.

OpenStudy (anonymous):

I'm just being asked to factor

OpenStudy (anonymous):

Ok then.

OpenStudy (anonymous):

what about 24t^3+3 is it the same concept?

OpenStudy (anonymous):

Essentially, yes. \[3(8t ^{3} +1)\] I could be mistaken though.

OpenStudy (radar):

This is the sum of two perfect cubes (and a factor 3) There is a formula for that a^3 + b^3=(a+b)(a^2-ab+b^2)

OpenStudy (radar):

3(2t+1)((2t)^2-2t+1)

OpenStudy (radar):

Might want to double check that.